2^3-10[4-(-2+18)3]
Answer by actuary(112)

The calculation is performed by following the "order of precedence of operations". First, do the calculaion inisde of the inner parenthesis, so -2+18=16. so the expression is now 2^3-10*(4-16*3)
Next, multiply 16 times 3 remembering that a negative sign is in fron of the 16.
The expression is now 2^3-10
Now do the calculation inside of the parenthesis, remembering the sign of the answer.
The expression is now 2^3-10*(-44)
Now multiply -10 times -44 (A negative number times a negative number is a positive number. So the expression is now 2^3+440
But 2^3=8
So we now have 8+440 and 8+440 = 448